OpenStudy (anonymous):

Its marketing people estimate that at a price of p dollars, it can sell a total of q = 260,000 − 10,000p copies. What price will bring in the greatest revenue? guys i need help asap my mom wants me to answer this problem before i go to bed plz!!!!!

OpenStudy (anonymous):

someone needs to help me

jimthompson5910 (jim_thompson5910):

Revenue = (# of items sold)*(price per item) Revenue = q*p R = (260,000 − 10,000p)*p R = 260,000p − 10,000p^2 The next step is to find the vertex of that quadratic. I'll let you take over.

OpenStudy (anonymous):

dude you gotta help me @jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you able to graph?

OpenStudy (anonymous):

yes but what am i suppose to graph?

jimthompson5910 (jim_thompson5910):

R = 260,000p − 10,000p^2 or y = 260,000x − 10,000x^2

OpenStudy (anonymous):

what does that ????? mean?

jimthompson5910 (jim_thompson5910):

refresh the page

jimthompson5910 (jim_thompson5910):

the symbols get weird, idk why

OpenStudy (anonymous):

is it 260000x times 10000x^2

jimthompson5910 (jim_thompson5910):

minus

jimthompson5910 (jim_thompson5910):

y = 260,000x - 10,000x^2 or y = -10,000x^2 + 260,000x

OpenStudy (anonymous):

but there is no vertex?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

there is, you just need to have the right window

OpenStudy (anonymous):

do i enter both equations?

jimthompson5910 (jim_thompson5910):

you can also use the formula x = -b/(2a) a = -10,000 and b = 260,000 in this case

OpenStudy (anonymous):

omg thank you so much

OpenStudy (anonymous):

Marketing estimates that at a price of p dollars, it can sell q = 300,000 − 15,000p copies, but each copy costs $4 to make. What price will give the greatest profit?

OpenStudy (anonymous):

@jim_thompson5910 can you give me another equation for this one